# Integrable Stochastic Processes Associated with the $D_{2}$ Algebra

### v1

@@ 5 Conclusion
In this work, we construct an integrable stochastic processes associated with the $D_{2}$ algebra, comprising four types of particles distinguished by their charge values and colors. We demonstrate that this model can be decomposed into two XXX spin chains (or SSEPs). This decomposition enables us to derive the exact solution of the model directly from the well-established solution of XXX spin chain. Unlike many other high-rank quantum integrable models, we do not need to solve the coupled Bethe ansatz equations for the rapidities [26, 27, 28]. This significantly simplifies the analytical solution of the system. Due to the decomposition property of the spectrum and the eigenstates, the observables can also be calculated analytically.
Our model can be also described as two independent lanes where particles are confined to hop within their respective lane, with inter-lane transitions prohibited. This constraint implies that the dynamics of the overall two-lane system decompose into the product of two independent, single-lane stochastic processes.
In this work, all bulk transition rates are symmetric. However, our recent studies indicate that the model can be generalized to asymmetric case.
Figure 8: Generalization of the stochastic process (its ASEP analogue).
In this case, the $R$-matrix factorizes into a product of two XXZ spin chain $R$-matrices
where
$\mathcal{R}^{\sigma}_{\bar{1},\bar{2}}(u)$ $=\left(\begin{array}[]{cccc}\sinh(u+\eta_{1})&0&0&0\\ 0&\sinh u&\sinh\eta_{1}&0\\ 0&\sinh\eta_{1}&\sinh u&0\\ 0&0&0&\sinh(u+\eta_{1})\end{array}\right),$ (5.6)
$\mathcal{R}^{\tau}_{\tilde{1},\tilde{2}}(u)$ $=\mathcal{R}^{\sigma}_{\bar{1},\bar{2}}(u)|_{\eta_{1}\to\eta_{2}}.$ (5.7)
As a consequence, the model decomposes into two ASEPs with completely independent hopping rates (see Figure 8). This model is thought to be related to the $D_{2}^{(1)}$ algebra [14, 29, 30, 31], and its detailed investigation will be the subject of our future research.
Another interesting question is the integrable stochastic process associated with the $D_{2}^{(2)}$ algebra [32, 33, 34, 35]. The model corresponding to the $D_{2}$ algebra can be decomposed into stochastic processes on two independent lanes, the $D_{2}^{(2)}$ algebra, however, can lead to stochastic processes on two coupled lanes, where particles can hop between two lanes. This suggests the existence of “two-dimensional” integrable stochastic processes. It has been addressed that the transfer matrix of the $D_{2}^{(2)}$ spin chain can be factorized as the product of transfer matrices of two staggered XXZ spin chains with fixed spectral difference after a global gauge transformation. In this case, the model more closely resembles a two-lane stochastic processes [36, 37, 38, 39, 40, 41].

### v2

@@ 6 Conclusion
In this work, we first construct integrable stochastic processes associated with the $D_{2}$ algebra, comprising four types of particles distinguished by their charge values and colors. We demonstrate that our model can be decomposed into two XXX spin chains (or SSEPs). This decomposition enables us to derive the exact solution of the model directly from the well-established solution of XXX spin chain. Unlike many other high-rank quantum integrable models, we do not need to solve the coupled Bethe ansatz equations for the rapidities [32, 33, 34]. This significantly simplifies the analytical solution of the system. Due to the decomposition property of the spectrum and the eigenstates, the observables can also be calculated analytically. We further examine the asymmetric analogue, which likewise possesses a factorized form and admits decomposition into two ASEPs.
The models presented in this paper can be described as two independent lanes where particles are confined to hop within their respective lane, with inter-lane transitions prohibited. This constraint implies that the dynamics of the overall two-lane system decompose into the product of two independent, single-lane stochastic processes.
One of our future objectives is to investigate the integrable stochastic process related to the $D_{2}^{(2)}$ algebra [35, 36, 37, 38]. The model corresponding to the $D_{2}$ algebra can be decomposed into stochastic processes on two independent lanes, the $D_{2}^{(2)}$ algebra, however, can lead to stochastic processes on two coupled lanes, where particles can hop between two lanes. This suggests the existence of “two-dimensional” integrable stochastic processes. It has been addressed that the transfer matrix of the $D_{2}^{(2)}$ spin chain can be factorized as the product of transfer matrices of two staggered XXZ spin chains with fixed spectral difference after a global gauge transformation. In this case, the model more closely resembles a two-lane stochastic processes [39, 40, 41, 42, 43, 44].